OpenStudy (anonymous):

differentiate: y=arctan(sqrt:(1-x/1+x))

zepdrix (zepdrix):

oo that's a doozy :U getting confused on the arctan derivative part of it or what?

zepdrix (zepdrix):

\[\Large\bf\sf y\quad=\quad \arctan\left(\sqrt{\frac{1-x}{1+x}}\right)\]

OpenStudy (anonymous):

haha yeah its scary

OpenStudy (nincompoop):

it is not scary

OpenStudy (anonymous):

yes thats correct @zepdrix

zepdrix (zepdrix):

\[\Large\bf\sf y=\arctan(\color{royalblue}{x})\]\[\Large\bf\sf y'=\frac{1}{1+(\color{royalblue}{x})^2}\]So we'll have to keep this derivative in mind.

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

\[\Large\bf\sf y=\arctan\left(\color{royalblue}{\sqrt{\frac{1-x}{1+x}}}\right)\] \[\Large\bf\sf y'=\frac{1}{1+\left(\color{royalblue}{\sqrt{\frac{1-x}{1+x}}}\right)^2}\cdot\color{orangered}{\left(\sqrt{\frac{1-x}{1+x}}\right)'}\]

zepdrix (zepdrix):

We've got a couple things going on here, the derivative of arctan is probably what you expect, right? The entire argument is stuffed into that square. But we also need to apply the chain rule. So we need to take the derivative of that orange thing.

OpenStudy (anonymous):

yes and would you then say that the square roots are changed to the 1/2 and -1/2?

OpenStudy (anonymous):

then quotient rule?

zepdrix (zepdrix):

yah derivative of square root, then chain into quotient rule. very good.

zepdrix (zepdrix):

\[\Large\bf\sf y'=\frac{1}{1+\left(\color{royalblue}{\sqrt{\frac{1-x}{1+x}}}\right)^2}\cdot\color{orangered}{\left(\frac{1}{2\sqrt{\dfrac{1-x}{1+x}}}\right)}\cdot\color{green}{\left(\dfrac{1-x}{1+x}\right)'}\]

zepdrix (zepdrix):

brb i need moar chocolate milk! :O

zepdrix (zepdrix):

mmm

OpenStudy (anonymous):

haha ok idk what ya did there in orange

zepdrix (zepdrix):

You can apply the power rule `if you want`. But that's a derivative that is really really worth memorizing. derivative of sqrt is `one over 2sqrts`.\[\Large\bf\sf \frac{d}{dx}\sqrt x\quad=\quad \frac{1}{2\sqrt x}\]

OpenStudy (anonymous):

oh never seen that one before, only the power rule! will write this one down!, so then we finally take the derivative of the last one in green which is -1?

zepdrix (zepdrix):

for your quotient rule? Hmm no you should** get something quite different from -1

zepdrix (zepdrix):

\[\Large\bf\sf \left(\frac{1-x}{1+x}\right)'\quad=\quad \frac{(1-x)'(1+x)-(1-x)(1+x)'}{(1+x)^2}\]

OpenStudy (anonymous):

-2x?

zepdrix (zepdrix):

\[\Large\bf\sf =\quad \frac{(-1)(1+x)-(1-x)(1)}{(1+x)^2}\] \[\Large\bf\sf =\quad \frac{-1-x-1+x}{(1+x)^2}\]

OpenStudy (anonymous):

oops forgot the bottom: -2x/(1+x^2)

OpenStudy (anonymous):

i mean the bottom is (1-x)^2

zepdrix (zepdrix):

ok good. And it looks like the x's cancel on top, yes? Leaving us with -2

zepdrix (zepdrix):

@Stew.a.r.t. I know this isn't exactly the same problem you're working on. But look at the steps at least. Might help you through yours.

zepdrix (zepdrix):

-x+x on top right? Make sure you distributed your negative signs correctly.

OpenStudy (anonymous):

umm i know im wrong but wouldit be -2x/(1+2x-x)?? sorry its been awhile since ive done algebra lol

zepdrix (zepdrix):

Don't expand out the square in the bottom. Leave it alone. I still don't understand how you're getting -2x on top. You should be getting -2.

zepdrix (zepdrix):

After taking derivative, your numerator should be, (-1)(1+x) - (1-x) did you end up with that?

OpenStudy (anonymous):

yes over (1+x)^2

zepdrix (zepdrix):

Distribute the first negative, -1-x - (1-x) distribute the `second negative` -1-x -1+x did you forget about the second negative?

OpenStudy (anonymous):

oh ok so the top would be -2 cause the x's cancel

zepdrix (zepdrix):

\[\Large\bf\sf y'=\frac{1}{1+\left(\color{royalblue}{\sqrt{\frac{1-x}{1+x}}}\right)^2}\cdot\color{orangered}{\left(\frac{1}{2\sqrt{\dfrac{1-x}{1+x}}}\right)}\cdot\color{green}{\dfrac{-2}{(1+x)^2}}\]Ok good! :)

zepdrix (zepdrix):

It can probably be simplified a little bit.. but uhhhh whatever. That's a fine answer.

OpenStudy (anonymous):

WOW thats crazy sorry it took me awhile lol...so you wouldn't have to simplify at all?

zepdrix (zepdrix):

I dunno, I'm not your teacher lol. Really depends on what your teacher wants. Like notice that we have a square root being squared in our tangent derivative. And we're diving by 2 and multiplying by 2 in a few different spots. It can be cleaned up a little bit. Nothing to fuss over though :o

OpenStudy (anonymous):

ok i dont think he would mind if i left it like this but just to clarify so the twos inthe two different spots could you cancel them into a 1 and -1?

OpenStudy (anonymous):

thanks a bunch!!! :))

zepdrix (zepdrix):

\c:/